decide whether the function is even, odd, or


neither.


f(x) = x^4+ 3

1
Expert's answer
2021-11-07T17:36:39-0500

Taking into account that the domain "\R" of the function "f(x) = x^4+ 3" is symmetric about the "y"-axis and "f(-x) = (-x)^4+ 3 = x^4+ 3=f(x)," we conclude that this function is even.